How many grams of calcium oxide (CaO) would be needed to neutralise 0.5 litre of battery acid (H2SO4) with a concentration of 0.25mol/L?
Given the atomic masses:
1
Expert's answer
2018-08-29T04:43:26-0400
Mr(CaO) = 56g/mol
m(CaO) = 0.5*0.25*56 = 7(g)